# SurgRIPE challenge: Benchmark of Surgical Robot Instrument Pose Estimation

### v1

@@ Abstract
Accurate instrument pose estimation is a crucial step towards the future of robotic surgery, enabling applications such as autonomous surgical task execution. Vision-based methods for surgical instrument pose estimation provide a practical approach to tool tracking, but they often require markers to be attached to the instruments. Recently, more research has focused on the development of marker-less methods based on deep learning. However, acquiring realistic surgical data, with ground truth instrument poses, required for deep learning training, is challenging. To address the issues in surgical instrument pose estimation, we introduce the Surgical Robot Instrument Pose Estimation (SurgRIPE) challenge, hosted at the 26th International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI) in 2023. The objectives of this challenge are: (1) to provide the surgical vision community with realistic surgical video data paired with ground truth instrument poses, and (2) to establish a benchmark for evaluating markerless pose estimation methods. The challenge led to the development of several novel algorithms that showcased improved accuracy and robustness over existing methods. The performance evaluation study on the SurgRIPE dataset highlights the potential of these advanced algorithms to be integrated into robotic surgery systems, paving the way for more precise and autonomous surgical procedures. The SurgRIPE challenge has successfully established a new benchmark for the field, encouraging further research and development in surgical robot instrument pose estimation.
keywords:
Surgical Instrument Pose Estimation , Instrument Tracking , Robot-assisted Minimally Invasive Surgery
@@ 1 Introduction
Robot-assisted Minimally Invasive Surgery (RMIS) has evolved significantly in the last decade driven by advances in artificial intelligence (AI) and surgical robotics. Platforms like the da Vinci®system have revolutionised surgical procedures by providing enhanced instrument control and intraoperative visualisation, greatly improving surgical assistance. Accurate pose estimation of surgical instruments has become a crucial task in RMIS, as it is essential for enabling applications such as autonomous surgical task execution Wang et al. (2018), surgical skill assessment Gao et al. (2014), and surgical workflow analysis Lecuyer et al. (2020).
@@
Commercial external devices, such as depth cameras and electromagnetic trackers Federico et al. (2019), can provide accurate instrument pose estimation. However, their applicability intraoperatively is limited due to space requirements and hardware setup constraints in the Operating Room. Another solution for surgical instrument pose estimation is the use of kinematic information Wang et al. (2022) from the integrated joint encoders of robotic platforms such as the da Vinci®system. Although it does not rely on extra hardware, this method requires additional hand-eye calibration and suffers from estimation errors due to the complexity of the cable-driven robotic system Cui et al. (2023).
@@
Marker-based vision methods for surgical instrument pose estimation use external markers to simplify the task Cartucho et al. (2021). They are limited by the fact that they rely on the marker always being visible in the camera’s Field-of-View (FOV) and are sensitive to background variations such as light reflection and occlusion. Moreover, these markers do not directly reflect the pose of the instrument, requiring the calculation of complex geometrical transformations. Therefore, markerless methods offer a promising and practical approach to surgical instrument tracking without hardware modifications.
@@
Object pose estimation has been well studied in the computer vision literature, with benchmarks like LineMODBrachmann et al. (2014) and YCB-VideoXiang et al. (2018) utilising RGBD sensors and ArUco markers for 6 Degrees of Freedom (DoF) pose estimation in non-medical scenes. These datasets have facilitated significant advancements for pose estimation methods for natural scene tasks. However, a comparable benchmark for 6DoF pose estimation for surgical tasks and environments is lacking. Existing medical datasets focus on the processing of 2D information as shown in Table. 2. These datasets neglect the 3D information that is required for the estimation of 6DoF pose. For example, EndoVis18 RobSeg Allan et al. (2020), EndoVis17 RobSeg Allan et al. (2019) and ROBUST-MIS Roß et al. (2021) provide data sets with 2D segmentation annotations. EndoVisPoseDu et al. (2018) provides data with ground truth 2D instrument keypoints. SuPerLi et al. (2019) provides ground truth kinematics information. However, in the latter case, the relevant 6DoF pose cannot be derived directly from the kinematics information.
@@
State-of-the-art methods, such as Peng et al. (2019)Wen et al. (2023), have been established for natural scene 6DoF object pose estimation tasks. However, due to the lack of surgical benchmarks and data sets, translating these methods to RMIS is difficult due to the following factors which are unique to surgery:
@@
    Partial object visibility. The limited operating space in RMIS means that the endoscopic camera remains very close to the surgical instruments, allowing only partial visibility within the camera’s field of view (FOV). This partial visibility hinders the performance of some state-of-the-art (SOTA) pose estimation methods due to the common requirements for full object visibility.
@@
    Surgical scene variations and occlusions. In RMIS, surgical tools interact with soft tissue and organs, leading to potential occlusions of the tool tip (e.g. due to blood), making pose estimation unstable. In addition, variations in the surgical scene, such as lighting conditions and specular reflections, further affect the accuracy of pose estimation.
@@
    High precision requirement. Pose estimation datasets often use RGBD cameras to generate ground truth data, resulting in errors of centimetre scale. However, given that the typical diameter of surgical tools is around 5 millimetres, the accuracy requirements in RMIS are of millimetre scale.
Figure 1: SurgRIPE data collection pipeline
To address the above issues, we present the SurgRIPE challenge, hosted at the 26th International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI) in 2023. This paper first introduces the SurgRIPE dataset, which has been created for markerless estimation of the 6DoF pose of the end-effector of surgical instruments. To acquire accurate and consistent ground truth surgical instrument poses while capturing video data in a realistic surgical setup, the novel pipeline shown in Fig. 1 was used which combines marker-based pose estimation with deep learning-based image inpainting. A keydot marker is used to get ground truth 6DoF pose data, which is then removed using a deep-learning inpainting model Suvorov et al. (2021) to avoid generating any shortcut visual cues which could bias the pose estimation. Finally, 3D models are used to generate segmentation masks of the surgical instruments.
@@
Table 1: Description of the cases on the training subset
Dataset Labelled Frames Type Annotation Type
SurgRIPE 2841 Real Endoscope 3D Pose & Segmentation Mask
ROBUST-MIS Roß et al. (2021) 10040 Real Endoscope Segmentation Mask
SuPer Li et al. (2019) 2000 Real Stereo Endoscope Kinematics Info & Segmentation Mask
EndiVisPose Du et al. (2018) 1850 Real Endoscope 2D keypoints & Segmentation Mask
EndoVis17 RobSeg Allan et al. (2019) 3000 Real Endoscope Segmentation Mask
EndoVis18 RobSeg Allan et al. (2020) 5700 Real Stereo Endoscope Segmentation Mask
Table 2: Comparison of the Surgical Tool Localization Dataset
2 Datasets and Annotation
2.1 Data
All SurgRIPE video data was captured using a da Vinci™Si endoscopic stereo camera, ensuring the acquisition of high-quality and clinically relevant images. Only images from the left camera were preserved and used for processing. Two different da Vinci surgical instruments were used, namely, the Large Needle Driver (LND) and the Maryland Bipolar Forceps (MBF). The collected videos include sequences without and with instrument occlusions. The sequences without occlusion included variations in lighting conditions and background scenes to diversify the data set and to simulate endoscopic environments such as when the light source is closer and farther from the tissue. The sequences with occlusion feature additional surgical tools, simulating scenarios with multiple instruments used by the surgeon.
The dataset is split into the LND and MBF subsets, with each subset containing only its respective instrument. Each subset contains the 3D model of the instrument, the pinhole camera intrinsic matrix, RGB images capturing the instrument moving over a surgical scene, segmentation masks of both the background and the instrument, and ground truth instrument poses. These poses correspond to the 6DoF movement of the wrist joint of the instruments, as shown in Fig. 2.
Figure 2: Projection of the 3d model onto the 2D image
The dataset information is illustrated in Table. 1. The LND dataset was extracted from 17 video clips containing 16737 raw frames and the MBF dataset was extracted from 15 video clips containing 11527 raw frames. All raw frames were captured at a framerate of 25 Hz. Finally, the LND subset consists of 1147 video frames without occlusion which are used for training, 373 frames without occlusion for testing, and 238 frames with occlusion for testing. The MBF subset contains 1,069 video frames without occlusion which are used for training, 209 frames without occlusion for testing, and 387 frames with occlusion for testing. In both cases, the occlusions were created using surgical instruments such as scissors and forceps. Each subset contains the following data.
Instrument Model The instrument 3D model is acquired from the da Vinci Research Kit (dVRK) wiki page Kazanzides et al. (2014) which contains 3D models of multiple Endowrist instruments. This challenge focuses on the joint part in the 3D model, as shown in Fig. 2.
@@ 2.1 Data
Segmentation Mask The segmentation mask of the instrument is generated by projecting the 3D model of the instrument onto the 2D image given the camera’s intrinsic parameters and the tool pose. In our case, only the joint part of each instrument is considered. Fig 6 illustrates how the segmentation map is generated from the pose and 3D model.
@@
Figure 3: Non-occlusion Image Sample
Figure 4: Occlusion Image Sample
@@ 2.2 Data annotation
To estimate the ground truth (GT) of the 6DoF pose of a surgical instrument, we designed a holder with a keydot marker Bradski (2000) which is attached to the joint of the instrument as shown in the raw image in Fig 1. The pose of the keydot marker $T_{M}^{C}$ is estimated via a PnP solver Lepetit et al. (2009). To recover the instrument pose $P^{I}$, the transformation from the keydot marker to the tip of the instrument $T_{I}^{M}$ needs to be estimated via the 3D pose alignment script as shown in Fig. 5.
Given the estimated marker pose and the camera intrinsic matrix, the 3D model of the instrument is projected onto the 2D image. Since the marker pose does not coincide with the instrument pose, the projection of the 3D model will not align with the target instrument area. To refine this 3D-to-2D projection, the Graphical User Interface (GUI) shown in Fig. 5 has been implemented in OpenCV to manually adjust the transformation from the marker to the instrument tip $T_{I}^{M}$ until the instrument model projection perfectly aligns with the target area. The adjustment of the parameters of the marker-instrument transformation is repeated until the 3D model projection overlays the instrument on every frame of the dataset. This results in a precise transformation between the marker and the instrument, and therefore in accurate and consistent GT instrument pose estimation. Compared with some previous work Allan et al. (2018) that acquired manual annotation for every frame, our pipeline can achieve abundant high-precision and consistent ground truth annotations with minimal risk of human error.
Figure 5: GUI for 3D Pose Alignment
Figure 6: Segmentation Generation Pipeline
@@ 2.2.1 Image inpainting
Since in a real surgical scenario, the keydot marker would not be present, we used image inpainting Suvorov et al. (2021) to remove the keydot marker from the images used to train and test the compared models, as shown in Fig. 1. When inpainting was applied, we used random size masks to ensure that the method could not generate any consistent visual features that can be learnt by any computer vision method. To our knowledge, our proposed method is the first work that can generate a large surgical instrument 6DoF pose dataset of realistic images with high-accuracy ground truth annotations which can be used for training and testing of deep learning models.
@@ 3.1 Performance evaluation protocol
Our benchmarking metrics include standard performance evaluation metrics (translation and rotation error) as well as the Benchmark for 6D Object Pose Estimation (BOP) metrics Brachmann et al. (2014), which define a protocol to evaluate the pose estimation accuracy. BOP metrics provide comprehensive evaluation metrics and have been widely applied in pose estimation tasks. The Average Accuracy metric is also proposed here. We have split the performance evaluation metrics into primary and secondary. The former category is used to rank the compared methods.
@@ 3.3 Secondary Metrics
2D projection metric (proj2d): This metric evaluates how well the 3D points of an object when projected into the 2D image plane using the predicted pose and camera intrinsic parameters, match the ground truth 2D projections. The aim is to measure the average distance between the corresponding projected points from the estimated pose and the ground truth pose. Given the camera intrinsic matrix $\mathbf{K}$, the ground truth pose $(\mathbf{R}_{gt},\mathbf{t}_{gt})$, and the predicted pose, $(\mathbf{R}_{pred},\mathbf{t}_{pred})$, the point cloud $\mathbf{X}$ can be projected to the 2D point sets $\mathbf{p}^{\text{gt}}$ and $\mathbf{p}^{\text{pred}}$, respectively. The proj2d error is computed as
@@ 3.4 Winner identification protocol
The teams were ranked according to their ADD and Ave Acc. The challenge winners are the submissions with the highest ADD and Avg Acc scores on the test dataset comprehensively.

### v2

@@ Abstract
Accurate instrument pose estimation is a crucial step towards the future of robotic surgery, enabling applications such as autonomous surgical task execution. Vision-based methods for surgical instrument pose estimation provide a practical approach to tool tracking, but they often require markers to be attached to the instruments. Recently, more research has focused on the development of markerless methods based on deep learning. However, acquiring realistic surgical data, with ground truth (GT) instrument poses, required for deep learning training, is challenging. To address the issues in surgical instrument pose estimation, we introduce the Surgical Robot Instrument Pose Estimation (SurgRIPE) challenge, hosted at the 26th International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI) in 2023. The objectives of this challenge are: (1) to provide the surgical vision community with realistic surgical video data paired with ground truth instrument poses, and (2) to establish a benchmark for evaluating markerless pose estimation methods. The challenge led to the development of several novel algorithms that showcased improved accuracy and robustness over existing methods. The performance evaluation study on the SurgRIPE dataset highlights the potential of these advanced algorithms to be integrated into robotic surgery systems, paving the way for more precise and autonomous surgical procedures. The SurgRIPE challenge has successfully established a new benchmark for the field, encouraging further research and development in surgical robot instrument pose estimation.
Keywords: Surgical Instrument Pose Estimation, Instrument Tracking, Robot-assisted Minimally Invasive Surgery
journal: Medical Image Analysis
@@ 1 Introduction
Robot-assisted Minimally Invasive Surgery (RAMIS) has evolved significantly in the last decade driven by advances in artificial intelligence (AI) and surgical robotics. Platforms like the da Vinci®system have revolutionised surgical procedures by providing enhanced instrument control and intraoperative visualisation, greatly improving surgical assistance. Accurate pose estimation of surgical instruments has become a crucial task in RAMIS, as it is essential for enabling applications such as autonomous surgical task execution (Wang et al., 2018), surgical skill assessment (Gao et al., 2014), and surgical workflow analysis (Lecuyer et al., 2020).
@@
Commercial external devices, such as depth cameras and electromagnetic trackers (Federico et al., 2019), can provide accurate instrument pose estimation. However, their applicability intraoperatively is limited due to space requirements and hardware setup constraints in the Operating Room. Another solution for surgical instrument pose estimation is the use of kinematic information (Wang et al., 2022) from the integrated joint encoders of robotic platforms such as the da Vinci®system. Although it does not rely on extra hardware, this method requires additional hand-eye calibration and suffers from estimation errors due to the complexity of the cable-driven robotic system (Cui et al., 2023).
@@
Marker-based vision methods for surgical instrument pose estimation use external markers to simplify the task (Cartucho et al., 2021). They are limited by the fact that they rely on the marker always being visible in the camera’s Field-of-View (FOV) and are sensitive to background variations such as light reflection and occlusion. Moreover, these markers do not directly reflect the pose of the instrument, requiring the calculation of complex geometrical transformations. Therefore, markerless methods offer a promising and practical approach to surgical instrument tracking without hardware modifications.
Dataset Labelled Frames Type Annotation Type
SurgRIPE 2841 Real Endoscope 3D Pose & Segmentation Mask
ROBUST-MIS (Roß et al., 2021) 10040 Real Endoscope Segmentation Mask
SuPer (Li et al., 2019) 2000 Real Stereo Endoscope Kinematics Info & Segmentation Mask
EndiVisPose (Du et al., 2018) 1850 Real Endoscope 2D keypoints & Segmentation Mask
EndoVis17 RobSeg (Allan et al., 2019) 3000 Real Endoscope Segmentation Mask
EndoVis18 RobSeg (Allan et al., 2020) 5700 Real Stereo Endoscope Segmentation Mask
Table 1: Comparison of the surgical tool localization datasets.
@@
Object pose estimation has been well studied in the computer vision literature, with benchmarks like LineMOD (Brachmann et al., 2014) and YCB-Video (Xiang et al., 2018) utilising RGBD sensors and ArUco markers for 6 Degrees of Freedom (DoF) pose estimation in non-medical scenes. These datasets have facilitated significant advancements in pose estimation methods for natural scene tasks. However, a comparable benchmark for 6DoF pose estimation for surgical tasks and environments is lacking. Existing medical datasets focus on the processing of 2D information as shown in Table 1. These datasets neglect the 3D information that is required for the estimation of 6DoF pose. For example, EndoVis18 RobSeg (Allan et al., 2020), EndoVis17 RobSeg (Allan et al., 2019) and ROBUST-MIS (Roß et al., 2021) provide datasets with 2D segmentation annotations. EndoVisPose (Du et al., 2018) provides data with ground truth 2D instrument keypoints. SuPer (Li et al., 2019) provides ground truth kinematic information. However, in the latter case, the relevant 6DoF pose cannot be derived directly from the kinematic information.
@@
State-of-the-art (SOTA) methods, such as (Peng et al., 2019; Wen et al., 2023), have been established for natural scene 6DoF object pose estimation tasks. However, due to the lack of surgical benchmarks and datasets, translating these methods to RAMIS is difficult due to the following factors which are unique to surgery:
@@
    Partial object visibility. The limited operating space in RAMIS means that the endoscopic camera remains very close to the surgical instruments, allowing only partial visibility within the camera’s field of view. This partial visibility hinders the performance of some state-of-the-art pose estimation methods due to the common requirements for full object visibility.
@@
    Surgical scene variations and occlusions. In RAMIS, surgical tools interact with soft tissue and organs, leading to potential occlusions of the tool tip (e.g. due to blood), making pose estimation unstable. In addition, variations in the surgical scene, such as lighting conditions and specular reflections, further affect the accuracy of pose estimation.
@@
    High precision requirement. Pose estimation datasets often use RGBD cameras to generate ground truth data, resulting in errors of centimetre scale. However, given that the typical diameter of surgical tools is around 5 millimetres, the accuracy requirements in RAMIS are of millimetre scale.
To address the above issues, we present the SurgRIPE challenge, hosted at the 26th International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI) in 2023. This paper first introduces the SurgRIPE dataset, which has been created for markerless estimation of the 6DoF pose of the wrist mechanism of surgical instruments. To acquire accurate and consistent ground truth surgical instrument poses while capturing video data in a realistic surgical setup, the novel pipeline shown in Fig. 1 was used which combines marker-based pose estimation with deep learning-based image inpainting. A keydot marker is used to get ground truth 6DoF pose data, which is then removed using a deep-learning inpainting model (Suvorov et al., 2021) to avoid generating any shortcut visual cues which could bias the pose estimation. Finally, 3D models are used to generate segmentation masks of the surgical instruments.
Figure 1: SurgRIPE data collection pipeline. The GT pose was captured with a keydot marker which was removed using image inpainting. The 3D model of the tool was used to generate the segmentation mask.
@@
Figure 2: Non-occlusion Image Sample.
Figure 3: Occlusion Image Sample.
2 Datasets and Annotation
2.1 Data
All SurgRIPE video data was captured using a da Vinci™Si endoscopic stereo camera, ensuring the acquisition of high-quality and clinically relevant images. All the data was acquired in the Hamlyn Centre, Imperial College London. Only images from the left camera were preserved and used for processing. Two different da Vinci surgical instruments were used, namely, the Large Needle Driver (LND) and the Maryland Bipolar Forceps (MBF). The collected videos include sequences without and with instrument occlusions. The sequences without occlusion included variations in lighting conditions and background scenes to diversify the dataset and to simulate endoscopic environments such as when the light source is closer and farther from the tissue. The sequences with occlusion feature additional surgical tools, simulating scenarios with multiple instruments used by the surgeon.
The dataset is split into the LND and MBF subsets, with each subset containing only its respective instrument. Each subset contains the 3D model of the instrument, the pinhole camera intrinsic matrix, RGB images capturing the instrument moving over a surgical scene, segmentation masks of both the background and the instrument, and ground truth instrument poses. These poses correspond to the 6DoF movement of the wrist joint of the instruments.
The dataset information is illustrated in Table 2. The LND dataset was extracted from 17 video clips containing 16747 raw frames and the MBF dataset was extracted from 15 video clips containing 11527 raw frames. All raw frames were captured at a framerate of 25 Hz. Finally, the LND subset consists of 1147 video frames without occlusion which are used for training, 373 frames without occlusion for testing, and 238 frames with occlusion for testing. The MBF subset contains 1,069 video frames without occlusion which are used for training, 209 frames without occlusion for testing, and 387 frames with occlusion for testing. In both cases, the occlusions were created using surgical instruments such as scissors and forceps.
@@ 2.1 Data
Table 2: Description of the generated dataset.
To enable the participating teams to evaluate their algorithms before submitting the final results, sample images from the test sets were provided before the final submission deadline. The sample LND test sets include 5 frames without occlusion and 3 frames with occlusion. The sample MBF test sets also include 5 frames without occlusion and 3 frames with occlusion.
Both LND and MBF subsets contain the following data:
Figure 4: (Left) The raw RGB image. (Middle) Projection of the 3D model onto the raw image given the estimated 6DoF pose. (Right) The final segmentation result.
Instrument Model The instrument 3D model is acquired from the da Vinci Research Kit (dVRK) wiki page (Kazanzides et al., 2014) which contains 3D models of multiple Endowrist instruments. This challenge focuses on the joint part in the 3D model, as shown in Fig. 5.
@@
Figure 5: Projection of the 3D model onto the 2D image.
Segmentation Mask The segmentation mask of the instrument is generated by projecting the 3D model of the instrument onto the 2D image given the camera’s intrinsic parameters and the tool pose. In our case, only the joint part of each instrument is considered. Fig. 4 illustrates how the segmentation map is generated from the pose and 3D model.
@@
Figure 6: GUI for 3D Pose Alignment.
@@ 2.2 Data annotation
To estimate the ground truth of the 6DoF pose of a surgical instrument, we designed a holder with a keydot marker (Bradski, 2000) which is attached to the joint of the instrument as shown in the raw image in Fig. 1. The pose of the keydot marker $T_{M}^{C}$ is estimated via a PnP solver (Lepetit et al., 2009). To recover the instrument pose $P^{I}$, the transformation from the keydot marker to the tip of the instrument $T_{I}^{M}$ needs to be estimated via the 3D pose alignment script as shown in Fig. 6.
Given the estimated marker pose and the camera intrinsic matrix, the 3D model of the instrument is projected onto the 2D image. Since the marker pose does not coincide with the instrument pose, the projection of the 3D model will not align with the target instrument area. To refine this 3D-to-2D projection, the Graphical User Interface (GUI) shown in Fig. 6 has been implemented in OpenCV to manually adjust the transformation from the marker to the instrument tip $T_{I}^{M}$ until the instrument model projection perfectly aligns with the target area. The adjustment of the parameters of the marker-instrument transformation is repeated until the 3D model projection overlays the instrument on every frame of the dataset. This results in a precise transformation between the marker and the instrument, and therefore in accurate and consistent GT instrument pose estimation. Compared with some previous work (Allan et al., 2018) that acquired manual annotation for every frame, our pipeline can achieve abundant high-precision and consistent ground truth annotations with minimal risk of human error.
@@ 2.2.1 Image inpainting
Since in a real surgical scenario, the keydot marker would not be present, we used image inpainting (Suvorov et al., 2021) to remove the keydot marker from the images used to train and test the compared models, as shown in Fig. 1. When inpainting was applied, we used random size masks to ensure that the method could not generate any consistent visual features that can be learnt by any computer vision method. To our knowledge, our proposed method is the first work that can generate a large surgical instrument 6DoF pose dataset of realistic images with high-accuracy ground truth annotations which can be used for training and testing of deep learning models.
@@ 3.1 Performance evaluation protocol
Our benchmarking metrics include standard performance evaluation metrics (translation and rotation error) as well as the Benchmark for 6D Object Pose Estimation (BOP) metrics (Brachmann et al., 2014), which define a protocol to evaluate the pose estimation accuracy. BOP metrics provide comprehensive evaluation metrics and have been widely applied in pose estimation tasks. The Average Accuracy metric is also proposed here. We have split the performance evaluation metrics into primary and secondary. The former category is used to rank the compared methods.
@@ 3.3 Secondary Metrics
2D projection metric (proj2d): This metric evaluates how well the 3D points of an object match their ground truth 2D projections. The 3D points are projected onto the 2D image plane using the predicted pose and camera intrinsic parameters. The aim is to measure the average distance between the corresponding projected points from the estimated pose and the ground truth pose. Given the camera intrinsic matrix $\mathbf{K}$, the ground truth pose $(\mathbf{R}_{gt},\mathbf{t}_{gt})$, and the predicted pose, $(\mathbf{R}_{pred},\mathbf{t}_{pred})$, the point cloud $\mathbf{X}$ can be projected to the 2D point sets $\mathbf{p}^{\text{gt}}$ and $\mathbf{p}^{\text{pred}}$, respectively. The proj2d error is computed as
@@ 3.4 Winner identification protocol
The teams were ranked according to their ADD and Avg Acc. The challenge winners are the submissions with the highest ADD and Avg Acc scores on the test dataset comprehensively.
4 Challenge organization
All submissions were uploaded to the Synapse platform using Docker containers. Participants’ code can be made available via emailing the corresponding team authors. The participants could submit multiple times before the submission deadline.
All supervised and semi-supervised methods were allowed in this challenge. Considering there were abundant surgical instrument datasets proposed before, the participants were encouraged to use existing datasets along with SurgRIPE training data during the training phase.
Data presented in the challenge can be used for publication purposes only after the first version of the joint publication summarizing challenge results is submitted. All the team participants could be qualified as authors. The participants can publish their own method after the first version of the joint publication. The dataset is available to participating teams and other interested parties under CC BY 4.0 (Attribution).
The timetable for the challenge is shown in Table 3. The training data (LND, MBF) was released on 2023 May 15th, followed by the sample test data on June 1st, allowing participants to familiarize themselves with the dataset and benchmark their solutions. The evaluation phase started on September 15th, during which participants were required to submit their solutions in the form of Docker containers. The final submission deadline was October 5th at 11:59 PM GMT. The challenge was hosted during the Endoscopic Vision Challenge workshop as part of MICCAI 2023 on October 8th, where results and insights were presented.
Dates Events
2023 May, 15th Release of training data (LND, MBF)
2023 June, 1st Release of sample test data
2023 September, 15th Start of evaluation
2023 October, 5th Submission deadline
2023 October, 8th Challenge and Representation Day
2024 December, 31th Acceptance of post-challenge submissions
Table 3: Timetable for the SurgRIPE challenge.
The first-place award was £1000 and the second-place award was £500. The awards and rankings were announced at the EndoVis 2023 challenge workshop. Late submissions are included in this report, however, not eligible for awards. Members of the organizers’ institutes could participate but were not eligible for awards.
The work presented in this paper follows the Biomedical Image Analysis Challenges (BIAS) Reporting Guideline (Maier-Hein et al., 2020) as required by MedIA.